OpenStudy (anonymous):

Find an equation of the line that goes through the point (2,1) and has a slope of 4. Leave equation in point-slope form

OpenStudy (anonymous):

We need a graph to figure this out.

OpenStudy (anonymous):

ok well your slope is 4 so your trying to get y=mx+b! so plug in your 4 into the m then your X value into the X and the Y and into the Y so it should be 1=4(2)+b then multiply 4 and 2 then subtract 8 and get b=-7 then put it back so it is y=4x-7

OpenStudy (anonymous):

I got 2-1= m(3-2) thats not right?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

point slope form is y-y1 = m(x-x1) In your case, m = 4, x1 = 2 and y1 = 1